OpenStudy (anonymous):

Simplify (4xy^2)^3(xy)^5 a.) 64x^8y^11 b.) 64x^15y^30 c.) 12x^2y^11 d.) 12x^8y^11

OpenStudy (anonymous):

i think its (a). (4xy^2)^3(xy)^5 = 64 (xy^2)^3(xy)^5 = 64 (x^3)(y^6)(x^5)(y^5) = 64x^8(y^11)

OpenStudy (anonymous):

Thanks(: